Given: C = 500 + 0.8Yd where Yd = Y – T G = 200 T...


Given:
C = 500 + 0.8Yd where Yd = Y – T
G = 200
T = 180
I = 100

Calculate the equilibrium level of national income.
Answer #1

Solution

At equilibrium,
Y = C + I + G
C = 500 + 0.8Yd, Yd = Y – T
I = 100
G = 200
T = 180
Now,
Yd = Y -T
Therefore,
Y = Yd + T
Y = Yd + 180
So, At equilibrium,
Yd + 180 = 500 + 0.8Yd + 100 + 200
1.8Yd = 800 – 180
Yd = 620/1.8
Yd = 344.44
Now we know, Y = Y -T

So,
344.44 = Y – 180
Y = 524.44
Therefore, the equilibrium level of national income is 524.44